# Think, Reflect, Create: Metacognitive Learning for Zero-Shot Robotic Planning with LLMs

### v1

@@ 5 Conclusions
In this paper, we introduced an early-stage framework to explore the research question: Can LLMs be empowered with metacognitive capabilities to reason, reflect, and create, thereby enhancing their ability to perform robotic tasks with minimal demonstrations? Our proposed framework integrates metacognitive learning into LLM-powered multi-robot collaboration, inspired by human reflective problem-solving processes. By equipping LLM-driven robotic agents with a metacognitive learning module, our framework enables effective reasoning, self-reflection on failures, and creative synthesis of novel solutions in zero-shot robotic planning scenarios. Experimental results demonstrate that our framework outperforms state-of-the-art methods and can generate innovative task solutions that differ from the provided ground-truth plans. These findings highlight the potential of metacognitive strategies to significantly enhance the adaptability, reliability, and creativity of robotic systems powered by LLMs.
Limitations
While this early-stage work shows exciting results on exploiting metacognitive learning methodology to advance LLM-based robot operations. Several limitations remain. For example, the current metacognitive self-reflection focuses primarily on identifying missing or misapplied skills but does not yet support finer-grained failure modes such as multi-agent coordination errors and long-horizon dependency mistakes. Extending the system to reason hierarchically about such failures remains an open challenge.

### v2

@@ Conclusions
This paper presented an initial effort in developing a metacognitive learning framework that integrates reasoning, self-reflection, and structured creativity into LLM-powered multi-robot collaboration. Specifically, our approach enables robotic agents to self-reflect on failures and synthesize novel solutions in zero-shot scenarios. We also introduced a new multi-robotic collaboration task and evaluated our method on both existing and new benchmark tasks. Experimental results show that our framework outperforms state-of-the-art methods and generates creative solutions beyond the ground-truth plans, demonstrating its effectiveness in enhancing the adaptability, reliability, and creativity of LLM-powered robotic systems.